A physical therapist wants to determine the difference in the proportion of men and women who participate in regular sustained physical activity. What sample size should be obtained if hehe wishes the estimate to be within fourfour percentage points with 9595​% ​confidence, assuming that
​(a) hehe uses the estimates of 21.321.3​% male and 18.318.3​% female from a previous​ year?
​(b) hehe does not use any prior​ estimates?
​(a) nequals= nothing ​(Round up to the nearest whole​ number.)
​(b) nequals= nothing ​(Round up to the nearest whole​ number.)
Answer #1

a.)E=\pm 4%=0.04

p1=0.213, p2=0.183

At 95% confidence level Critical value=1.96 (use standard Normal table)

Sample Size:

n1 = n2 = [p1(1-pi) + p2(1-p2)ler

=[0.213(1-0.213)+0.183(0.183)](\frac{1.96}{0.04})^2

=761.4579

=761

b.)

he does not use any prior​ estimates

n1=n2=0.5*(\frac{z}{E})^2

=0.5*(\frac{1.96}{0.04})^2

=1200.5

1201